You purchase 170 shares for $60 a share ($10,200), and after a year the price falls to $55. Calculate the percentage return on your investment if you bought the stock on margin and the margin requirement was (ignore commissions, dividends, and interest expense):

15 percent. Use a minus sign to enter the amount as a negative value.

Round your answer to one decimal place. % 70 percent.

Use a minus sign to enter the amount as a negative value. Round your answer to one decimal place. % 85 percent.

Use a minus sign to enter the amount as a negative value. Round your answer to one decimal place. %
Answer #1

Investment amount

=10200*15%

=1530

gain/loss on investment=(55-60)*170=-850

the percentage return on your investment

=-850/1530

=-55.6%

the above is answer..